### v1

# {textblock*}

Abstract
In light of the latest Planck-Atacama Cosmology Telescope (P-ACT) joint results on the primordial scalar power spectrum, we show that the $R^{2}$ inflation model extended with a non-minimally coupled scalar field $\chi$—namely the $R^{2}$-Higgs model—can yield a larger spectral index $n_{s}$ and a small positive running $\alpha_{s}$ at cosmic microwave background (CMB) scales, which are consistent with the data. This is because the $\chi$ field contributes a blue-tilted component to the primordial power spectrum which both modifies the large-scale power and, as a result, significantly enhances power on small scales. The consequent enhancement can be large enough to lead to the formation of primordial black holes (PBHs) of mass $\lesssim 10^{20}\mathrm{g}$ as a dark matter candidate. Furthermore, future observations of the small-scale power spectrum, CMB spectral distortions, and stochastic gravitational waves will provide decisive tests of this model and its predictions for PBHs.
preprint: APS/123-QED
I Introduction
The $R^{2}$ inflation [1, 2], also known as the Starobinsky inflation, has been long regarded as one of the most successful inflationary models, both theoretically well-motivated and perfectly fits the observations. Its predictions for the spectral index $n_{s}$ of the primordial power spectrum and the tensor to scalar ratio $r$ are in excellent agreement with the Planck 2018 [3] and BICEP/Keck observations. The model also gives a precise prediction of a negative running of spectral index $\alpha_{s}$, which is also consistent with the Planck results  [^1] According to the Planck 2018 result, $n_{s}=0.6949\pm 0.0042$ (68% CL),$\alpha_{s}\equiv dn_{s}/d\ln k=-0.0041\pm 0.0067$ (68% CL). The tensor to scalar ratio is constrained to $r<0.032$ (95% CL) by joint Planck and BICEP/Keck observations [88]..
However, a newly released result on the primordial power spectrum from the Atacama Cosmology Telescope (ACT) DR6 [5, 6] reports a slightly larger spectral index. Combined with Planck data at large-to-intermediate scales (denoted as P-ACT) and together with CMB lensing and BAO data from DESI [7] (denoted as P-ACT-LB), the analysis yields a new constraint on the spectral index: $n_{s}=0.9743\pm 0.0034$ (68% CL) at the CMB pivot scale $k_{\mathrm{CMB}}=0.05~{}\mathrm{Mpc}^{-1}$ [5]. This result appears to disfavor the $R^{2}$ model as illustrated in Fig. 4. In addition, the error bar for the running of the spectral index $\alpha_{s}$ is tightened in the P-ACT-LB result, which suggests a slight positive running of the spectral index $\alpha_{s}=0.0062\pm 0.0052$ (68% CL) [6]. The value is not consistent with most of the single-field models that involve a convex potential, including the $R^{2}$ model. Several works [8, 9, 10, 11, 12, 13, 14, 15, 16] discuss extensions to the single-field $R^{2}$ model that may resolve this inconsistency. In this letter, we present a simple extension to the $R^{2}$ model by introducing a non-minimally coupled scalar field $\chi$ (also called $R^{2}$-Higgs model [17, 10, 18, 19, 11]) that may provide a better fit to the newly released data, thus potentially ”save” the $R^{2}$ model via a hybrid-like inflation scenario. The $\chi$ field, acting as a waterfall field, introduces a blue-tilted contribution to the total primordial power spectrum, thereby leading to a positive shift in both the spectral index $n_{s}$ and its running $\alpha_{s}$. Notably, this blue-tilted contribution peaked at small scales can trigger the collapse of overdense regions into primordial black holes (PBHs) [20, 21, 22, 23, 24] during radiation domination, which could constitute cold dark matter.
Therefore, we also discuss the possible hint from the P-ACT results to the small-scale features in the primordial power spectrum. In particular, we find that the scenarios involving the formation of primordial black holes with mass below $10^{20}\mathrm{g}$ through the enhanced small-scale power spectrum, accounting for all of the dark matter, show good agreement with the latest data. The growth of the power spectrum involved in our model also gives rise to distinctive, potentially detectable signatures in both the $\mu$-distortion and the stochastic gravitational wave background, which can serve as distinguishing features in future surveys like PIXIE [25, 26], LISA [27], and etc.
II The $R^{2}$-Higgs model
We consider the action where the $R^{2}$ gravity is non-minimally coupled by a scalar field $\chi$ [17, 10, 18, 19] with a slightly broken $Z_{2}$ symmetry characterized by $\chi_{0}$  [^2] The $Z_{2}$ breaking term is introduced in hybrid models to resolve the overproduction of domain wall and avoid the unacceptably huge quantum fluctuations [89, 19],
$S_{J}=\int d^{4}x\sqrt{-g}\ \bigg{[}\frac{M_{\mathrm{pl}}^{2}}{2}f(R,\chi)-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-V(\chi)\bigg{]},$ (1)
where $M_{\mathrm{pl}}\equiv(8\pi G)^{-1/2}$, $f(R,\chi)$ is given by  [^3] See an extension to $R^{3}$ in [11]
$f(R,\chi)=R+\frac{R^{2}}{6M^{2}}-\frac{\xi R}{M_{\mathrm{pl}}^{2}}(\chi-\chi_{0})^{2},$
and $V(\chi)$ is a quartic potential,
$V(\chi)=\frac{1}{4}\lambda(\chi^{2}-m^{2})^{2}.$
Performing conformal transformation $\tilde{g}_{\mu\nu}=Fg_{\mu\nu}$, the action (1) can be transformed into the Einstein frame [30] as
$S_{\mathrm{E}}=\int d^{4}x\sqrt{-\tilde{g}}\left[\frac{M_{\mathrm{pl}}^{2}}{2}\tilde{R}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right.$ (2)
$\left.-\frac{1}{2}F^{-1}\tilde{g}^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi-U(\phi,\chi)\right],$
where the $R^{2}$ term converted into a scalar degree of freedom $\phi$, given by
$\phi$ $\equiv M_{\mathrm{pl}}\sqrt{\frac{3}{2}}\ln F,$
and with an effective potential written as
$U(\phi,\chi)\equiv\frac{3}{4}M^{2}M_{\mathrm{pl}}^{2}W^{2}(\phi,\chi)+\frac{V(\chi)}{F(\phi)^{2}},$
$W(\phi,\chi)\equiv 1-\frac{1}{F(\phi)}\left[1-\xi\left(\frac{\chi-\chi_{0}}{M_{\mathrm{pl}}}\right)^{2}\right].$
In this framework, a hybrid-like inflation could be realized in the composition of two separated stages of effective single-field inflation dominated individually by $\phi$ and $\chi$ and an intermediate transition stage, which are described as follows:
  • St-1: Effective $R^{2}$ inflation. Initially, the inflaton slow-rolls from the plateau of the potential along the $\phi$ direction. Given the large value of $\phi/M_{\mathrm{pl}}$, the spectator field $\chi$ acquires an effective mass approximately given by $m_{\chi}^{2}=3M^{2}\xi+\mathcal{O}(F^{-1})$.
  • Transition. Effective single-field $R^{2}$ inflation ends when the inflaton reaches the potential minimum in the $\phi$ direction. Simply put, the universe experienced a short but complete version of $R^{2}$ inflation before this moment of transition. During this phase, $F$ approaches unity, $\phi$ becomes heavy ($m_{\phi}^{2}>9H^{2}/4$), and the inflaton dynamics begin to shift toward the $\chi$ direction.
  • St-2: Effective Higgs inflation. The slow roll along $\chi$ starts, with the heavy field $\phi$ stabilized in its potential minimum. The inflation will be ended in effective single-field configuration.
As we explained above, the scalaron $\chi$, acting as a spectator field (or equivalently, the isocurvature component) during the effective $R^{2}$ inflation, dominates the inflation when the St-2 starts. Such background evolution, commonly observed in hybrid inflation models [31, 32, 33] with an initially heavy isocurvature field $\chi$, can induce a sharp $k^{3}$ enhancement in the comoving curvature power spectrum and thus lead to PBH production.
A similar scenario arises in the curvaton model [34, 35, 36, 37, 38]. Unlike the waterfall field in hybrid inflation, the curvaton remains effectively massless and nearly static during inflation. It dominates the universe once the Hubble parameter becomes smaller than the curvaton mass after the end of inflation. In such cases, the comoving curvature power spectrum contributed by curvaton field perturbations that decays later should be close to scale-invariant, depending on the curvaton’s effective mass during inflation.
Notably, both cases are encompassed within the rich phenomenology of the $R^{2}$–Higgs inflationary model. During the first stage of inflation, the effective mass of $\chi$ is given by $m_{\chi}^{2}=F\partial^{2}U/\partial\chi^{2}\sim 3M^{2}\xi$. Since the Hubble parameter in this early stage is approximately $\sim M^{2}/4$, the ratio of $m_{\chi}^{2}$ to $H^{2}$ turns out to solely depend on the coupling constant $\xi$. Therefore, in principle, the model allows the power spectrum to grow from scale-invariant ($k^{0}$) to steeply blue-tilted ($k^{3}$) by varying the coupling constant $\xi$ from 00 to $\gtrsim 3/16$ [19]. Consequently, if a significant enhancement in the power spectrum occurs at small scales, a steeper growth helps to localize the feature, keeping it well separated from large-scale modes. On the other hand, when the spectrum grows more gently, the enhancement at small scales may reshaped the large-scale spectrum.
In the following, we present the $\xi$-dependence of the primordial power spectrum from the perspective of $\delta N$ formalism.
III Analytical power spectrum
The $\delta N$ formalism relates the comoving curvature perturbation $\mathcal{R}$ to the field perturbations, which are evaluated on spatially flat slicing at superhorizon scales [39, 40, 41, 42]. Here, $N$ denotes the number of e-folds from the flat slicing to the comoving slicing at the end of inflation, and $\delta N$ represents the perturbation in $N$ induced by the superhorizon field fluctuations.
For simplicity, we assume that the transition occurs instantaneously. In this case, for modes that exit the horizon before the transition, the total $\delta N$ measured at the end of inflation can be written as
$\mathcal{R}=\delta N=\delta N_{1}+\delta N_{2},$ (3)
[…]
Figure 1: A schematic diagram of the simplified trajectory. We assume that the transition stage occurs instantaneously at $n_{*}$, therefore the background evolution consists of two separate phases of effective single-field inflation along $\phi$ and $\chi$ directions, respectively. The fiducial background trajectory is indicated by black solid arrows, while the perturbed trajectory is shown with red solid arrows. For the perturbed trajectory, the boundary condition perturbed by the field perturbations $\delta\phi_{k}$ (dashed blue arrow) and $\delta\chi_{k}$ (dashed purple arrows) in the spatially flat gauge at the horizon exiting stage of mode $k$, denoted by $n_{k}$.
If we conduct the $\delta N$ calculation nonlinearly, we could observe a large deviation of the probability distribution function of $\mathcal{R}$ from Gaussianity. Such non-Gaussian features are potentially detectable by upcoming observations. We leave this for future work.
In this work, we restrict the calculation of $\delta N$ to linear level [^4] See a more detailed analysis in [19]. Assuming $\delta\chi_{k}$ and $\delta\phi$ are small perturbations, and treating both St-1 and St-2 as effective single-field inflation, we obtain the expression for $\delta N$:
$\delta N=\left(\frac{\partial N_{1}}{\partial\phi}\delta\phi_{k}\right)_{n_{k}}+\left(\frac{\partial N_{2}}{\partial\chi}\delta\chi_{k}\right)_{n_{\star}},$ (4)
where the first term is equivalent to the $\delta N$ in $R^{2}$ inflation, and the extra contribution of $\delta\chi$ appears in the second term. Intuitively speaking, the second term, namely the $\delta N_{2}$, reflects the perturbation of the number of e-folds during the second stage of inflation, induced by the mode $\delta\chi_{k}$ that exits the horizon during the first stage. This term is expected to be exponentially suppressed for large-scale modes since the amplitude of $\delta\chi_{k}$ evolves as
$\delta\chi_{k}(n_{\star})\approx\delta\chi_{k}(n_{k})e^{-a(n_{\star}-n_{k})/2},$ (5)
according to the Mukhanov-Sasaki equation. The decay rate $a$ is given by
$a$ $\equiv\mathrm{Re}\left(3-3\sqrt{1-\frac{16m^{2}_{\chi}}{9M^{2}}}\right)$ (6)
$=\mathrm{Re}\left(3-3\sqrt{1-\frac{16}{3}\xi}\right),$
during the St-1, which equals the maximum $3$ for the heavy $\chi$ case and smaller when decreasing $\xi$.
Figure 2: The $n_{s}-\alpha_{s}$ relation shown in Eq. (8) for different $M_{\mathrm{PBH}}$s together with the joint constrains from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). To plot the figure, we use $\mathcal{P}_{\mathcal{R}}^{\mathrm{max}}/\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=10^{7}$ in order to have $f_{\mathrm{PBH}}=1$. The hollow black dots refer to cases that $a=3$, or equivalently $\xi\geq 3/16$ when the small-scale amplification is well isolated from the large scales.
As a result, for modes that exit the horizon before the transition stage, the power spectrum contains both contributions from the dominating light field $\phi$ as well as the isocurvature field $\chi$. Assuming a constant $H$ and Bunch-Davis vacuum deep inside of the horizon, the power spectrum for modes $k\leq k_{1}$, where $k_{1}$ is the mode that crosses the horizon during the end of the first phase of inflation (which is approximately the peak scale of the power spectrum), reads
$\mathcal{P}_{\mathcal{R}}(k)=\mathcal{P}_{\mathcal{R}}^{R^{2}}(k)+\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}\left(\frac{k}{k_{1}}\right)^{a}.$ (7)
Here the $\mathcal{P}_{\mathcal{R}}^{R^{2}}$ refers to the single field $R^{2}$ model power spectrum, and the peak value of the spectrum $\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}$ is a simple function of $\chi_{0}$ and $\lambda$. Accordingly, the spectral index, the running of the spectral index, and the tensor-to-scalar ratio at the CMB scale are given by
$n_{s}\approx 1-\frac{2}{N_{1}}+\frac{\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}e^{-aN_{1}}\left(\frac{2}{N_{1}}+a\right),$ (8)
$\alpha_{s}\approx-\frac{2}{N_{1}^{2}}+\frac{\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}e^{-aN_{1}}\left(\frac{2}{N_{1}^{2}}+\frac{2a}{N_{1}}+a^{2}\right),$
$r\approx\frac{2}{N_{1}^{2}},$
where $N_{1}$ is the number of e-folds before the end of the first stage of inflation corresponding to the CMB pivot scale. Since $k_{1}$ and $N_{1}$ correspond to the peak scale of the power spectrum, considering sharp-peak case, they are related to the PBH mass by
$M_{\mathrm{PBH}}\sim 3\times 10^{48}\mathrm{g}\times e^{-2N_{1}}.$ (9)
Therefore, to have PBH as all of the dark matter in the present universe, i.e. $f_{\mathrm{PBH}}=1$ is not yet constraint from observation, $N_{1}=30\sim 40$ is needed  [^5] See reviews on the observational constraints on PBH as dark matter [45, 47, 76].
When fixing the ratio of the power spectrum value at the peak scale and the CMB scale to be ${\mathcal{P}_{\mathcal{R}}^{\mathrm{peak}}}/{\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}}\sim 10^{7}$ (in order to generate a sizable abundance of primordial black holes today, i.e. $f_{\mathrm{PBH}}\sim 1$ [45, 46, 47]), if the coupling constant $\xi$ is larger than $3/16$ ($a=3$), the exponential suppression renders the last term in $n_{s}$ and $\alpha_{s}$ negligible, which means that the small scale can remain localized without affecting the large scale. In this limit, the large-scale power spectrum reduces to that of the single-field $R^{2}$ inflation, with $\alpha_{s}\lesssim-0.0013$, $n_{s}\lesssim 0.95$ for $N_{1}\lesssim 40$. When $e^{-aN_{1}}$ is of order $10^{-7}$ or less, the modification to the large-scale spectrum comes from the small-scale enhancement can be significant.
Case 1 2 3
Parameters
$M/M_{\rm pl}$ $1.35\times 10^{-5}$ $\ 1.78\times 10^{-5}$ $3.5\times 10^{-5}$
$m/M$ $1/6$ $\ 1/6$ $1/6$
$\xi$ $0.0530$ $0.0689$ $0.1163$
$\chi_{0}/M$ $0.28$ $0.39$ $0.39$
CMB Observables
$n_{s}(0.05\mathrm{Mpc}^{-1})$ $0.975$ $0.973$ $0.973$
$n_{s}(0.002\mathrm{Mpc}^{-1})$ $0.966$ $0.955$ $0.912$
$\alpha_{s}(0.05\mathrm{Mpc}^{-1})$ $0.008$ $0.014$ 0.0550.0550.0550.055666Excluded at more than $2\sigma$ level.
PBHs
$M_{\mathrm{PBH}}/\mathrm{g}$ $1.5\times 10^{8}$ $6\times 10^{18}$ $1.7\times 10^{26}$
$f_{\mathrm{PBH}}$ $\sim 1$ $\sim 1$ $\ll 1$
Table 1: The benchmark parameters. The primordial power spectrum is normalized by the Planck 2018 result $\mathcal{P}_{\mathcal{R}}^{\mathrm{CMB}}=\mathcal{P}_{\mathcal{R}}(0.05\mathrm{Mpc}^{-1})=2.1\times 10^{-9}$. A joint P-ACT-LB analysis reports a spectral index $n_{s}(0.05{\rm Mpc}^{-1})=0.9743\pm 0.0034$. The total number of efolds $N$ of the inflation in our model can be tuned from $50\sim 60$ without affecting the IR by adjusting $\lambda$, thus not listed in the table.
Figure 3: The primordial power spectrum measured at the end of inflation together with the (future) observational constraints. The black dashed line shows the result for $\xi>3/16$, whose power spectrum converges to the effective single-field $R^{2}$ inflation. The curves are obtained by numerically solving for the primordial power spectrum by the Transport method using the model parameters from Table 1 case 2 except for a varying $\xi$. From the darkest to the lightest solid curves, we take $\xi=0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (case 2), $0.0685$. The constraint plot comes from [48], including CMB constraints (red) from Planck [3], Lyman-$\alpha$ forest constraints [49] (blue), $\mu$-distortion constraints (green) from FIRAS [50], and PIXIE (future) [25, 26], and gravitational wave constraints (purple) from PTAs [51, 52, 53], SKA [54] (future), LISA [27] (future) and BBO [55] (future).
IV Numerical results
To accurately account for the subtleties in the model, we carry out numerical computations of the perturbation evolution based on the Transport method [56]. We consider three sets of benchmark parameters, as listed in Table 1: Case 1 corresponds to ultra-light PBHs, Case 2 to asteroid PBHs within the dark matter window, and Case 3 to PBHs detectable via microlensing. All three parameter sets yield slightly larger spectral indexes at the CMB pivot scale of $0.05\mathrm{Mpc}^{-1}$ , which remain consistent with the P-ACT results.
However, according to P-ACT-LB constraint on $\alpha_{s}$, Case 3 is excluded at more than the $2\sigma$ level. In fact, in our model, the production of PBHs with masses $\gtrsim 10^{20}\mathrm{g}$ is generally ruled out by ACT, as it is associated with a relatively large running of the spectral index (see Fig. 2). As a result, we conclude that PBHs larger than $\sim 10^{20}$ g are excluded as a dark matter candidate at 2 $\sigma$ level according to the P-ACT results.
Case 1 can be both consistent with the data but also potentially interesting. The ultra-light PBHs produced are expected to evaporate before Big Bang Nucleosynthesis (BBN) due to Hawking radiation [57, 58, 59]. Those evaporating tiny PBHs can help to reheat the universe, produce high frequency gravitational waves by Hawking radiation, and induce second order gravitational waves through the poltergeist mechanism[60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. From quantum gravity point of view, those evaporated PBHs may abundantly leave remnants as dark matter [72, 73, 74, 75].
In the following, we focus on the most compelling case, Case 2, where the near monochromatic asteroid mass PBHs produced from the model can constitute the whole dark matter [76, 77, 78], which shows great compatibility with ACT results.
The full power spectra resulting from parameter Case 2 with varying $\xi$s are shown in Fig. 3. The variation of the spectral index and its running can be observed in Fig. 3, and more intuitively in Fig. 4 and Fig. 5, where the tendencies are consistent with the analytical approximations given by Eq. (8).
It is also worth mentioning that, as illustrated in Fig. 3, the enhancement of the small-scale power spectrum not only affects the large-scale spectral index $n_{s}$ and its running $\alpha_{s}$, which can be probed through direct measurements of the primordial power spectrum from future CMB surveys [79, 80], but also induces significant $\mu$-distortion [81, 82]—potentially detectable by PIXIE [25, 26] [^7] See detailed analysis in [11].. Additionally, the associated induced gravitational waves fall within the sensitivity ranges of future detectors such as LISA [27], DECIGO [84], Taiji [85]/TianQin [86] and BBO [55]. Therefore, upcoming missions could provide crucial observational evidence in support of this model.
Figure 4: The $n_{s}-r$ relation in the $R^{2}$ model and $R^{2}$-Higgs model together with the joint constraints from BICEP/Keck observations [87] (denoted as BK18) with Planck-LB (orange), ACT-LB (blue), and P-ACT-LB (purple). The constraint contours are taken from [6]. The dark blue dashed line with hollow dots refers to the $n_{s}-r$ relation in the $R^{2}$ model, and the light cyan dashed line and dots refer to the $R^{2}$-Higgs model with PBH production. For the light cyan plot, we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), $0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (star-shaped, Case 2), $0.0685$.
Figure 5: The $n_{s}-\alpha_{s}$ relation in $R^{2}$-Higgs model together with the joint constraints from Planck, P-ACT and P-ACT-LB (The constraint contours are taken from [6]). The light cyan dots refer to the $\alpha_{s}-n_{s}$ relation in $R^{2}$-Higgs model with PBH production, for which we use the model parameters from Table 1 Case 2 except for a varying $\xi$. From smaller $n_{s}$ (left) to larger $n_{s}$ (right), we take $\xi=0.3125$ (hollow dot), $0.0725$, $0.0715$, $0.0705$, $0.0695$, $0.0689$ (star-shaped, Case 2), $0.0685$.
V Discussion and Conclusion
The recently released ACT observation joint with Planck data reports a slightly larger spectral index $n_{s}$ and a positive running $\alpha$ placing well-motivated $R^{2}$ inflation and Higgs inflation models under tension at $2\sigma$ level. In this letter, we have shown that we can resolve the tension within the simple combination of the two models, namely, introducing a non-minimally coupled scalar field to the $R^{2}$ model. By tuning the non-minimally coupling constant $\xi$, the model realizes a hybrid-like inflationary scenario, in which the non-minimally coupled scalar field contributes a blue-tilted component to the primordial power spectrum, yielding an increase in the $n_{s}$ and a positive $\alpha_{s}$ consistent with the current observation.
At the same time, the blue-tilted spectrum significantly enhances the fluctuation power at small scales, potentially leading to the formation of primordial black holes (PBHs). In Fig. 2, we show the relation between $n_{s}$, $\alpha_{s}$, and the mass of the resulting PBHs. To remain consistent with ACT results, We demonstrated that the observed values of $n_{s}$ and positive $\alpha_{s}$ may serve as indirect evidence of such small-scale enhancement, consistent with the formation of PBHs with masses $\lesssim 10^{20}~{}\mathrm{g}$, which may account for the entirety of dark matter. It is worth mentioning that we do not claim that the tension between the new data and $R^{2}$ model predictions must be resolved by PBHs that explain all dark matter, but rather demonstrate our model naturally aligns the modifications of large scale power spectrum with testable small-scale features in the model. More specifically, the small-scale power spectrum of our model can be tested by upcoming observations of $\mu$-distortion (e.g. via PIXIE) and the stochastic gravitational wave background (e.g. via LISA), offering promising avenues for observational confirmation.
Acknowledgements.
We thank Misao Sasaki, Tsutomu Yanagida and Ying-li Zhang for useful discussions. This work is supported in part by JSPS KAKENHI Grant Nos. JP20H05853, JP24K00624 [X.W.]; JP23KF0289, JP24H01825, JP24K07027 [K.K.], and by Forefront Physics and Mathematics Program to Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) Program, the University of Tokyo [X.W.]. Kavli IPMU is supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. X.W. would also like to acknowledge the insightful input from the Random Discussion, held weekly at Kavli IPMU within the theoretical cosmology group.